a commercial bank will loan you 54567 for 4 years to buy a car. the loan must be repaid in equal monthly payments at the end of the month . the annual interest rate on the loan is 5-98% of the unpaid balance. what is the amount of the monthly payments?
Answer #1

Assuming the interest rate mentioned is 5.98%

Now, for this question, present value of the monthly payments made for 4 years at 5.98% (yearly rate). PV of an annuity is mathematically represented as:

r = 5.98%/12 = 0.01495% (monthly), n = 4 * 12 = 48 months

54567 = P * [\frac{1 - (1 + 0.0001495)^{-48}}{0.0001495}]

P = $1,140.98 --> Monthly payment